OpenStudy (anonymous):

Determine whether the lines through the pairs of points are parallel, perpendicular, or neither. (1, -3) and (3, -2); (0, -2) and (2, -3)

OpenStudy (anonymous):

find slopes of the lines between points: (1,-3) & (3,-2): m1=(-2-(-3))/(3-1) = 1/2 (0,-2) & (2,-3): m2=(-3-(-2))/(2-0)=-1/2 if lines parallel - should have the same slope; not the case=no parallel if lines perpendicular: m1=-1/m2; not the case= not perpendicular; so, the answer is = neither

OpenStudy (anonymous):

Thank you Thank you!!!